OpenStudy (anonymous):

Why does 5 * (13 choose 5) represent: the number of ways you can choose a team of five people chosen from a room of 13 people, where you are designating the captain? I understand the choose part, but not the 5 multiplier.

OpenStudy (unklerhaukus):

there can be 5 different captains

OpenStudy (anonymous):

I know that, but in the next "blank" why would it then be 13 choose 5? That isn't the number of people available remaining that we are allowed to pick.

OpenStudy (anonymous):

I really hope that makes sense.

OpenStudy (anonymous):

Shouldn't we only pick 4 after?

OpenStudy (unklerhaukus):

\[^{13}C_5\] is the number of possible teams, \[5\times~^{13}C_5\] is the number of possible teams where the captain is also chosen

OpenStudy (anonymous):

Oh, so there are 6 people all together?

OpenStudy (unklerhaukus):

no the captain is one of the 5 member in the team

OpenStudy (unklerhaukus):

members*

OpenStudy (anonymous):

Are we not already picking 5 people with \[\left(\begin{matrix}13 \\5 \end{matrix}\right)\]

OpenStudy (unklerhaukus):

yeah,

OpenStudy (anonymous):

So if we've chosen our 5, how would multiplying by 5 affect it?

OpenStudy (anonymous):

This is tricky.

OpenStudy (unklerhaukus):

im not sure what part you arn't getting @DanielHendrycks ?

OpenStudy (anonymous):

I now see. We are picking from the 5 people one captain. I thought we were picking from the entire pool a captain. There are 5 possibilities for captain and multiplying the two gives the numbers of varieties we can have of them (together).

OpenStudy (unklerhaukus):

thats right

OpenStudy (unklerhaukus):

\[^{13}C_5\times ^5C_1\]